In 2018, Laureen is currently single. She paid $2,640 of qualified tuition and related expenses for each of her twin daughters Sheri and Meri to attend State University as freshmen ($2,640 each for a total of $5,280). Sheri and Meri qualify as Laureen’s dependents. Laureen also paid $1,870 for her son Ryan’s (also Laureen’s dependent) tuition and related expenses to attend his junior year at State University. Finally, Laureen paid $1,370 for herself to attend seminars at a community college to help her improve her job skills.

What is the maximum amount of education credits Laureen can claim for these expenditures in each of the following alternative scenarios? (Leave no answer blank. Enter zero if applicable.)

a. Laureen's AGI is $45,000.

American opportunity credit ______

Lifetime learning credit ______


b. Laureen’s AGI is $95,000.     

American opportunity credit ______

Lifetime learning credit ______

c. Laureen’s AGI is $45,000 and Laureen paid $12,340 (not $1,870) for Ryan to attend graduate school (i.e, his fifth year, not his junior year).

American opportunity credit ______

Lifetime learning credit ______
Answer #1

Answer:-

a) Laureen's AGI - $45,000

For 2 daughter - AOTC is - (2000*2child)+(640*25%+2child)

=4000+320

=4320

For Ryan - 1870

AOTC - 6190

Laureen lifetime learning credit - Eligible is 2000 (The amount of the credit is 20 percent of the first $10,000 of qualified education expenses or a maximum of $2,000 per return)

so in above case it is - 1370*20% =274 (Since AGI is below clip of 52800 he can claim same)

=6190+274 = 6464 is eligible deduction

b)

Since AGI is 95000

AOTC can't be calimed if AGI is above 90000 and hence AOTC is zero and Lifetime learning credit can't be claimed if AGI is above 52800.. Hence it is zero education credit

c)

For Daughter it is same as a above i.e. 4,320

For Ryan it is = 2000+(10000*25%)or maximum 4000

=2000+2500 or 4000

so 4000 is allowed

so AOTC total of 8320 and LLC of 274 so claimed is 8,594

Hence Proved